OpenStudy (exoexoexo):

it took matt twice the time to drive from home to work as it took him to drive back home from work by the same route. if matt averaged 30 miles per hour(mph) on his drive to work, his average speed for the entire round trip was?

OpenStudy (anonymous):

ok so if it was 2 times as fast what would his drive backs mph be?

OpenStudy (exoexoexo):

60?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so since you have both of them how do you find the average

OpenStudy (exoexoexo):

(30+60)/2?

OpenStudy (anonymous):

yea so 45 would be it :)

OpenStudy (exoexoexo):

but its not an answer choice

OpenStudy (anonymous):

what are the choices???

OpenStudy (anonymous):

lol its not twice as fast its twice as slow

OpenStudy (exoexoexo):

30 25 40 15

OpenStudy (anonymous):

so (15+30)/2

OpenStudy (anonymous):

wait a minute

OpenStudy (anonymous):

@dan815 am i right in saying (15+30)/2 sorry im tired

OpenStudy (anonymous):

lol wait i think im just being dumb

OpenStudy (anonymous):

no wait im not it would be 22.5 mph are you sure thats the whole question

OpenStudy (exoexoexo):

I think it is a 30+2(15)/2

OpenStudy (anonymous):

no then that would just be 30

OpenStudy (exoexoexo):

30 mph is the average on Matt's drive to work while 15 mph is the average of Matt's drive to home from work

OpenStudy (anonymous):

because he took twice as long using the same rounte meaning he was twice as slow (15 mph) and the other was 30 so the avarage of that is 22.5mph

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so basicly in the context of the problem x= matts time from work so (x+(x/2))/2= avg

OpenStudy (anonymous):

and x=30 so

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

cool @EXOEXOEXO ???

OpenStudy (anonymous):

although its not a choice it is the right answer for what you gave me

OpenStudy (exoexoexo):

lol

OpenStudy (anonymous):

haha yea unless you made a typing error??

OpenStudy (exoexoexo):

I don't think I made one

OpenStudy (whpalmer4):

How long does it take him to drive to work? How long does it take him to drive home? Add them together to get the round trip time. Average speed for the round trip is round trip distance divided by round trip time.

OpenStudy (anonymous):

yea but thats not a choice :(

OpenStudy (whpalmer4):

That's all the information you need to calculate the answer... Let the distance to work be \(x\). Time to drive to work is \(x/30\). Time to drive home is \((1/2)*(x/30)\) Total distance driven is \(2x\) Average speed for trip is total distance driven divided by total time.

OpenStudy (anonymous):

yea i know that and you would get 22.5 but thats not an answer choice for him

OpenStudy (whpalmer4):

No, 22.5 is not the result of that calculation, correctly done.

OpenStudy (whpalmer4):

Time to drive to work is \[\frac{x}{30}\] Time to drive home is\[\frac{1}{2}*\frac{x}{30} = \frac{x}{60}\] Total time driving is \[\frac{x}{30}+\frac{x}{60} = \frac{3x}{60} = \frac{x}{20}\] average speed is \[\frac{2x}{\frac{x}{20}}=2x*\frac{20}{x} = \]

OpenStudy (whpalmer4):

Or, pick a distance, any distance, and work out the problem. Let's say it is 60 miles to work, to make the arithmetic easy. 60 miles at 30 miles per hour = 2 hours to drive to work Driving home takes 1/2 the time, so that's 1 hour to drive home. Total distance driven is 60+60 = 120 miles. Total time driving is 2+1 hours = 3 hours Average speed =